OpenStudy (anonymous):

can i get some help with finding the first 5 terms of taylor series of sin(2x) centered @ 0

OpenStudy (accidentalaichan):

Won't be giving the answers, but hopefully this will help you find them. http://www.math.ufl.edu/~vatter/teaching/m8w10/m8l12.pdf

OpenStudy (anonymous):

\[f(x)=2^0\sin2x\\ f'(x)=2^1\cos 2x\\ f''(x)=-2^2\sin2x\\ f'''(x)=-2^3\cos2x\\ ~~~~~~~~~~~\vdots\] For \(x=0\), all the even order derivatives will be zero, so let's just consider the odd (\(n=2k+1\)) powered ones. \[f^{(2k+1)}(x)=(-1)^{k}2^{2k+1}\cos2x~~~~~~~~~~\text{where }k=0,1,2,...\] To see why this formula works, you can plug in values of \(k\) and see if it matches up: \[k=0~\Rightarrow~f'(x)=2\cos2x\\ k=1~\Rightarrow~f'''(x)=-2^3\cos2x\\ \text{and so on...}\] So, the formula for a Taylor series centered at 0 is given by \[f(x)=\sum_{n=0}^\infty \frac{f^{(n)}(0)}{n!}x^n\] Or, in terms of odd values of \(n\), like above: \[f(x)=\sum_{k=0}^\infty \frac{f^{(2k+1)}(0)}{(2k+1)!}x^{2k+1}\] Noting that \(f^{(2k+1)}(0)=(-1)^{k}2^{2k+1}\), you have \[f(x)=\sum_{k=0}^\infty \frac{(-1)^{k}2^{2k+1}}{(2k+1)!}x^{2k+1}\] Now you have a formula you can use to find the first five terms.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome